OpenStudy (anonymous):

A particle of mass 4m explodes into 3 pieces of masses m,m,2m. The equal masses move along X & Y axes with velocities 4m/s & 6m/s respectively. Find the magnitude of velocity the heavier mass (Choices --Under root 17,2 x under root 13,under root 13 or under root 13/2 m/s

OpenStudy (anonymous):

 since the momentum of the system was zero at the beginning after the explode it remains zreo. now we have for the two masses : speed = sqrt(6^2+4^2) = sqrt(52) = 2sqrt(13) so the magnitude of the momentum of the two masses is m*2sqrt(13) this should be equal to 2m * |V| where |V| is the magnitude of the velocity of the larger mass 2m*|V| = m * 2sqrt(13) |V| = sqrt(13)